# Quantifying and Alleviating Co-Adaptation in Sparse-View 3D Gaussian Splatting

### v2

@@ 1 Introduction
Recent advances in 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) have demonstrated remarkable capabilities for photorealistic novel view synthesis in dense-view settings. However, under sparse-view supervision, 3DGS often suffers significant performance degradation in novel view rendering, manifesting as artifacts caused by incorrect geometry or appearance distribution due to limited supervision from training views. Existing works mainly improve sparse-view 3DGS through geometry regularization, such as monocular depth constraints (Li et al., 2024; Zhu et al., 2024; Xu et al., 2024; Xiong, 2024; Zhang et al., 2025a; Paliwal et al., 2024; Zheng et al., 2025; Chung et al., 2024), matching-based point initialization (Han et al., ), and dense point initialization (Wang et al., 2024; Fan et al., 2024). Although these methods achieve improvements by enhancing geometric accuracy, few have investigated appearance artifacts in sparse-view 3DGS. These artifacts, often overlooked, are common in novel view rendering and typically manifest as occasional color outliers—colors not belonging to the scene—as illustrated in Figure 1. In this paper, we focus on analyzing and mitigating appearance artifacts in the novel view rendering of sparse-view 3DGS.
@@
We begin our analysis by revisiting the rendering characteristics of 3DGS. In the splatting process, each pixel receives contributions from multiple Gaussians projected onto its corresponding tile on the image plane Kerbl et al. (2023). This multi-Gaussian composition is a fundamental characteristic of 3DGS, enabling effective appearance modeling under dense-view settings. However, it also introduces potential entanglement between Gaussians, especially when they are jointly optimized to fit overlapping pixel regions in the training views. In sparse-view scenarios, such entanglement becomes problematic—multiple Gaussians with divergent colors may aggressively collaborate to fit limited training pixels, ignoring the true underlying appearance distribution. Here, we refer to this entanglement as co-adaptation in 3DGS, and hypothesize that appearance artifacts in sparse-view novel view synthesis are closely related to excessive co-adaptation.
@@ 2 Related Work
3D Gaussian Splatting under Sparse Views. Recent studies have explored a variety of strategies to improve 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) under sparse-view supervision. Some approaches (Li et al., 2024; Zhu et al., 2024; Xu et al., 2024; Xiong, 2024; Zhang et al., 2025a; Paliwal et al., 2024; Zheng et al., 2025; Chung et al., 2024) incorporate external geometry priors, typically from monocular depth estimators (Birkl et al., 2023; Yang et al., 2024), to enforce multi-view depth consistency and enhance geometric fidelity. Other works (Liu et al., 2024a; Wu et al., 2025; Zhong et al., 2025; Yu et al., 2024; Liu et al., 2024b; Sun et al., 2024; Gao et al., 2024) leverage generative priors from diffusion models to synthesize plausible unseen views or guide view-dependent effects. In addition, other methods (Han et al., ; Zhang et al., 2024; Radl et al., 2024) introduce self-supervised or physically motivated regularization strategies, such as opacity decay, binocular-guided photometric consistency, and dual-model co-regularization, to suppress floating artifacts and improve geometric consistency. Additionally, a separate line of work (Xu et al., 2024; Zhang et al., 2025a; Charatan et al., 2024; Chen et al., 2024a, b; Ye et al., 2024a; Chen et al., 2024c; Zhang et al., 2025b; Ziwen et al., 2024) departs from per-scene optimization by training neural predictors across large-scale data to directly infer 3DGS representations from sparse inputs. Recent works (Xu et al., 2025; Park et al., 2025) have also explored dropout-based strategies for improving sparse-view 3DGS performance. While both of them report clear improvements, they attribute the effectiveness of dropout to empirical factors—such as reducing overfitting through fewer active splats (Xu et al., 2025), or enhancing gradient flow to distant Gaussians (Park et al., 2025). In contrast, our work identifies and formalizes co-adaptation suppression as the key underlying mechanism, offering a more principled explanation for why dropout benefits sparse-view generalization in 3DGS.
@@
Co-Adaptation in Neural Networks. Co-adaptation describes the tendency of neurons in neural networks to become overly dependent during training, leading to overfitting and poor generalization. This issue was first highlighted by Hinton et al. (2012), who proposed Dropout to mitigate such dependency by randomly deactivating neurons. Subsequent stochastic regularization methods (Wan et al., 2013; Gastaldi, 2017; Huang et al., 2016; Krueger et al., 2016; Wu et al., 2021; Verma et al., 2019; Hendrycks et al., 2019) further promote feature diversity through randomization, consistency enforcement, or latent-space mixing. Beyond empirical methods, theoretical studies (Baldi and Sadowski, 2013; Novak et al., 2018; Gal and Ghahramani, 2016) have analyzed co-adaptation’s negative impact on generalization, linking it to sharp loss minima (Baldi and Sadowski, 2013), low sensitivity to perturbations (Novak et al., 2018), and poor uncertainty estimation (Gal and Ghahramani, 2016). While these insights have been validated in classification and regression tasks, their implications for structured prediction like 3D reconstruction remain underexplored. Our work extends this investigation to 3D Gaussian Splatting, revealing similar interdependencies among spatial Gaussians and showing that dropout and opacity noise effectively mitigate co-adaptation under sparse-view supervision.
@@ 3 Co-Adaptation in Gaussian Splatting
In this chapter, we investigate co-adaptation in 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023), which significantly affects generalization under sparse-view settings. First, we explain co-adaptation in 3DGS and analyze why excessive co-adaptation leads to artifacts in novel view synthesis. Then, we propose a metric to quantify co-adaptation via variance across dropout-rendered outputs. Based on this metric, we make empirical observations about co-adaptation in 3DGS. Finally, we introduce two simple strategies to explicitly mitigate co-adaptation: dropout regularization and opacity noise injection.
@@ 3.1 Characterizing Co-Adaptation
Unlike explicit spatial representations (Gross and Pfister, 2011; Kaufman et al., 1993; Lorensen and Cline, 1998) that define fixed scene geometry, 3DGS represents scenes as a set of view-dependent Gaussians without explicit surface definitions. This allows 3DGS to model flexible, appearance-driven representations that adapt to different viewpoints. Specifically, 3DGS renders images by projecting 3D Gaussians onto the image plane and blending their contributions using differentiable alpha compositing. The color of each pixel $u$ is computed as:
@@ 3.2 Quantifying Co-Adaptation
Figure 4: Comparison of co-adaptation strength (CA), GS count and reconstruction quality under varying numbers of training views. (a-b) CA measured as pixel-wise variance for DNGaussian and Binocular3DGS on three target view types: extrapolation (far), interpolation (near), and training. (c) Gaussian count (left axis, in thousands) and average PSNR (right axis, in dB) for both methods. All plots are shown as functions of input view counts, based on the LLFF “flower” scene.
@@
    Increased training views reduce co-adaptation. Figure 4 shows that co-adaptation scores drop consistently as the number of training views increases, for both DNGaussian (Li et al., 2024) and Binocular3DGS (Han et al., ). While DNGaussian uses random initialization with a relatively stable Gaussian count in LLFF scenes (Mildenhall et al., 2019), Binocular3DGS leverages a pre-trained keypoint matcher (Truong et al., 2023) to generate denser initializations with more views. Despite these differences, both methods demonstrate that increasing view counts weakens co-adaptation and improves generalization.
@@ 3.3 Strategies for Alleviating Co-Adaptation
We explore two regularization strategies to mitigate excessive co-adaptation in 3D Gaussian Splatting. First, we adapt dropout (Hinton et al., 2012; Srivastava et al., 2014) to randomly mask subsets of Gaussians during rendering, reducing over-reliance on fixed combinations. Second, inspired by noise-based regularization (Bishop, 1995), we inject Gaussian noise into opacity parameters, encouraging more flexible Gaussian contributions. Both techniques aim to enhance generalization by loosening overly rigid dependencies between Gaussians.
@@ 4 Experiments
Method Setting LLFF (Mildenhall et al., 2019) DTU (Jensen et al., 2014)
@@ 4.1 Experiments Setup
Datasets. We conduct experiments on three datasets: LLFF (Mildenhall et al., 2019), DTU (Jensen et al., 2014), and Blender dataset (Mildenhall et al., 2021). Following prior works (Han et al., ; Niemeyer et al., 2022; Yang et al., 2023; Jain et al., 2021), we use 3 training views for LLFF and DTU, and 8 views for Blender. Test views follow prior settings (Han et al., ; Niemeyer et al., 2022; Yang et al., 2023; Jain et al., 2021). Input images are downsampled by a factor of 8 for LLFF, 4 for DTU, and 2 for Blender, relative to their original resolutions.
Baselines. We conduct experiments on 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023), DNGaussian (Li et al., 2024), FSGS (Zhu et al., 2024), CoR-GS (Zhang et al., 2024), and Binocular3DGS (Han et al., ). Our two proposed strategies, dropout regularization and opacity noise injection, are applied to each method to evaluate their effectiveness. Ablation studies are primarily conducted on Binocular3DGS. We follow the official implementations for all baselines. On LLFF, our reproduced results for DNGaussian and CoR-GS slightly differ from the reported values. For Blender, DNGaussian uses different training settings for different scenes, while we adopt a unified setup for fair comparison, which leads to lower results. We also re-run Binocular3DGS with white backgrounds to align with other methods. Baseline results in Table 2 are reported as original / reproduced, showing the original paper values and our re-implementation under the unified setup.
Metrics. To evaluate the rendering quality under novel views, we compute standard image quality metrics including average PSNR (peak signal-to-noise ratio), SSIM (Wang et al., 2004), and LPIPS (Zhang et al., 2018) on each dataset. To assess the impact of different strategies on co-adaptation, we additionally report Co-Adaptation Score (CA) (Section 3.2) computed on both training views and novel test views.
@@ 5 Conclusion
In this paper, we introduce the concept of co-adaptation in 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) and reveal its link to appearance artifacts under sparse-view settings. We propose the Co-Adaptation Score (CA) metric to quantify this effect and show that higher view density naturally reduces co-adaptation. Motivated by this, we propose two simple strategies—dropout regularization and opacity noise injection—that effectively mitigate co-adaptation and improve sparse-view novel view synthesis. We hope our findings inspire future work on mitigating co-adaptation in learned 3D representations.
@@ A.2 Extended Qualitative Results
We present additional qualitative results on the LLFF [Mildenhall et al., 2019], DTU [Jensen et al., 2014], and Blender [Mildenhall et al., 2021] datasets to further illustrate the visual improvements brought by our co-adaptation suppression strategies.
LLFF. We first examine the LLFF dataset, where improvements are clearly observed both in rendered images and Gaussian point clouds. As shown in Figure 9, our methods yield point clouds that are cleaner and more geometrically coherent. The baseline 3D Gaussian Splatting (3DGS) [Kerbl et al., 2023] often produces scattered, colorful speckles, which are significantly suppressed after applying our co-adaptation regularization. These visual improvements in point clouds directly lead to better rendering in novel views. For instance, in the trex scene, fossil bones exhibit high-frequency flickering and speckling in the baseline, which are clearly alleviated with our methods. In the flower scene, petals and leaves display incorrect colorization under novel views, while the regularized versions better preserve the natural appearance. Similarly, in the fern scene, severe color inconsistencies appear on the pole, ground, and leaf textures under the baseline, but these are significantly reduced when co-adaptation is suppressed. Overall, the results suggest that reducing co-adaptation enhances both structural consistency and appearance fidelity in sparse-view 3DGS.
@@
Method Comparison. Figure 12 presents a visual comparison across multiple 3DGS methods, using Binocular3DGS [Han et al., ] as the baseline. Binocular3DGS is a strong baseline with geometry-aware initialization, yet our co-adaptation suppression strategies (dropout and opacity noise) still lead to noticeable improvements. In the flower scene, existing sparse-view approaches exhibit occasional color speckles and geometry gaps; applying our methods further reduces these artifacts and yields more complete, coherent novel view reconstructions.
@@ A.3 Exploration of Noise Injection on Different Gaussian Attributes
where $s$ denotes the original scale and $\sigma$ is the noise strength. The results of this experiment are summarized in Table 6. Compared to the Binocular3DGS baseline, scale noise yields limited improvements in overall reconstruction metrics such as PSNR, SSIM [Wang et al., 2004], and LPIPS [Zhang et al., 2018]. However, it reduces the co-adaptation strength (CA) on both training and novel views, indicating that scale perturbation helps decouple Gaussian interactions to some extent. This finding aligns with our visual observations in Figure 13, where scale noise improves scene-level geometric coherence in some cases, but introduces mild local blurring in object details. The observed trade-off can be explained by the functional role of scale in Gaussian rendering: perturbing scale slightly alters the spatial support of Gaussians, which may improve coverage and fusion in sparse-view settings, but risks oversmoothing fine structures. As a result, while scale noise contributes to better scene fusion in some examples, it reduces local detail sharpness, leading to modest overall gains.
@@ A.7 Comparison with Alternative Strategies for Reducing Co-Adaptation
  • Baseline w/ AbsGS [Ye et al., 2024b]: We apply pixel-wise absolute gradient accumulation for Gaussian densification, replacing the standard gradient used in vanilla 3DGS. This method promotes the preservation of fine structures and reduces appearance artifacts, which may help alleviate view-specific co-adaptation.
  • Baseline w/o Opacity Decay [Han et al., ]: We remove the default multiplicative decay applied to Gaussian opacity parameters ($\times 0.995$ per iteration). Without this decay, redundant Gaussian opacities tend to maintain, causing higher co-adaptation scores, reduced sparsity, and more pronounced floating artifacts.
@@
w/ AbsGS [Ye et al., 2024b] 21.61 0.754 0.163 0.001446 0.001593
w/o opacity decay [Han et al., ] 20.26 0.708 0.202 0.004413 0.004830

### v3

@@ 1 Introduction
Recent advances in 3D Gaussian Splatting (3DGS) [1] have demonstrated remarkable capabilities for photorealistic novel view synthesis in dense-view settings. However, under sparse-view supervision, 3DGS often suffers significant performance degradation in novel view rendering, manifesting as artifacts caused by incorrect geometry or appearance distribution due to limited supervision from training views. Existing works mainly improve sparse-view 3DGS through geometry regularization, such as monocular depth constraints [2, 3, 4, 5, 6, 7, 8, 9], matching-based point initialization [10], and dense point initialization [11, 12]. Although these methods achieve improvements by enhancing geometric accuracy, few have investigated appearance artifacts in sparse-view 3DGS. These artifacts, often overlooked, are common in novel view rendering and typically manifest as occasional color outliers—colors not belonging to the scene—as illustrated in Figure 1. In this paper, we focus on analyzing and mitigating appearance artifacts in the novel view rendering of sparse-view 3DGS.
@@
We begin our analysis by revisiting the rendering characteristics of 3DGS. In the splatting process, each pixel receives contributions from multiple Gaussians projected onto its corresponding tile on the image plane [1]. This multi-Gaussian composition is a fundamental characteristic of 3DGS, enabling effective appearance modeling under dense-view settings. However, it also introduces potential entanglement between Gaussians, especially when they are jointly optimized to fit overlapping pixel regions in the training views. In sparse-view scenarios, such entanglement becomes problematic—multiple Gaussians with divergent colors may aggressively collaborate to fit limited training pixels, ignoring the true underlying appearance distribution. Here, we refer to this entanglement as co-adaptation in 3DGS, and hypothesize that appearance artifacts in sparse-view novel view synthesis are closely related to excessive co-adaptation.
@@ 2 Related Work
3D Gaussian Splatting under Sparse Views. Recent studies have explored a variety of strategies to improve 3D Gaussian Splatting (3DGS) [1] under sparse-view supervision. Some approaches [2, 3, 4, 5, 6, 7, 8, 9] incorporate external geometry priors, typically from monocular depth estimators [13, 14], to enforce multi-view depth consistency and enhance geometric fidelity. Other works [15, 16, 17, 18, 19, 20, 21] leverage generative priors from diffusion models to synthesize plausible unseen views or guide view-dependent effects. In addition, other methods [10, 22, 23] introduce self-supervised or physically motivated regularization strategies, such as opacity decay, binocular-guided photometric consistency, and dual-model co-regularization, to suppress floating artifacts and improve geometric consistency. Additionally, a separate line of work [4, 6, 24, 25, 26, 27, 28, 29, 30] departs from per-scene optimization by training neural predictors across large-scale data to directly infer 3DGS representations from sparse inputs. Recent works [31, 32] have also explored dropout-based strategies for improving sparse-view 3DGS performance. While both of them report clear improvements, they attribute the effectiveness of dropout to empirical factors—such as reducing overfitting through fewer active splats [31], or enhancing gradient flow to distant Gaussians [32]. In contrast, our work identifies and formalizes co-adaptation suppression as the key underlying mechanism, offering a more principled explanation for why dropout benefits sparse-view generalization in 3DGS.
@@
Co-Adaptation in Neural Networks. Co-adaptation describes the tendency of neurons in neural networks to become overly dependent during training, leading to overfitting and poor generalization. This issue was first highlighted by Hinton et al. [33], who proposed Dropout to mitigate such dependency by randomly deactivating neurons. Subsequent stochastic regularization methods [34, 35, 36, 37, 38, 39, 40] further promote feature diversity through randomization, consistency enforcement, or latent-space mixing. Beyond empirical methods, theoretical studies [41, 42, 43] have analyzed co-adaptation’s negative impact on generalization, linking it to sharp loss minima [41], low sensitivity to perturbations [42], and poor uncertainty estimation [43]. While these insights have been validated in classification and regression tasks, their implications for structured prediction like 3D reconstruction remain underexplored. Our work extends this investigation to 3D Gaussian Splatting, revealing similar interdependencies among spatial Gaussians and showing that dropout and opacity noise effectively mitigate co-adaptation under sparse-view supervision.
@@ 3 Co-Adaptation in Gaussian Splatting
In this chapter, we investigate co-adaptation in 3D Gaussian Splatting (3DGS) [1], which significantly affects generalization under sparse-view settings. First, we explain co-adaptation in 3DGS and analyze why excessive co-adaptation leads to artifacts in novel view synthesis. Then, we propose a metric to quantify co-adaptation via variance across dropout-rendered outputs. Based on this metric, we make empirical observations about co-adaptation in 3DGS. Finally, we introduce two simple strategies to explicitly mitigate co-adaptation: dropout regularization and opacity noise injection.
@@ 3.1 Characterizing Co-Adaptation
Unlike explicit spatial representations [44, 45, 46] that define fixed scene geometry, 3DGS represents scenes as a set of view-dependent Gaussians without explicit surface definitions. This allows 3DGS to model flexible, appearance-driven representations that adapt to different viewpoints. Specifically, 3DGS renders images by projecting 3D Gaussians onto the image plane and blending their contributions using differentiable alpha compositing. The color of each pixel $u$ is computed as:
@@ 3.2 Quantifying Co-Adaptation
Figure 4: Comparison of co‑adaptation strength (CA), GS count and reconstruction quality under varying numbers of training views. (a-b) CA measured as pixel-wise variance for DNGaussian and Binocular3DGS on three target view types: extrapolation (far), interpolation (near), and training. (c) Gaussian count (left axis, in thousands) and average PSNR (right axis, in dB) for both methods. All plots are shown as functions of input view counts, based on the LLFF “flower” scene.
@@
    Increased training views reduce co-adaptation. Figure 4 shows that co-adaptation scores drop consistently as the number of training views increases, for both DNGaussian [2] and Binocular3DGS [10]. While DNGaussian uses random initialization with a relatively stable Gaussian count in LLFF scenes [47], Binocular3DGS leverages a pre-trained keypoint matcher [48] to generate denser initializations with more views. Despite these differences, both methods demonstrate that increasing view counts weakens co-adaptation and improves generalization.
@@ 3.3 Strategies for Alleviating Co-Adaptation
We explore two regularization strategies to mitigate excessive co-adaptation in 3D Gaussian Splatting. First, we adapt dropout [33, 49] to randomly mask subsets of Gaussians during rendering, reducing over-reliance on fixed combinations. Second, inspired by noise-based regularization [50], we inject Gaussian noise into opacity parameters, encouraging more flexible Gaussian contributions. Both techniques aim to enhance generalization by loosening overly rigid dependencies between Gaussians.
@@ 4 Experiments
Method Setting LLFF [47] DTU [51]
@@ 4.1 Experiments Setup
Datasets. We conduct experiments on three datasets: LLFF [47], DTU [51], and Blender dataset [52]. Following prior works [10, 53, 54, 55], we use 3 training views for LLFF and DTU, and 8 views for Blender. Test views follow prior settings [10, 53, 54, 55]. Input images are downsampled by a factor of 8 for LLFF, 4 for DTU, and 2 for Blender, relative to their original resolutions.
Baselines. We conduct experiments on 3D Gaussian Splatting (3DGS) [1], DNGaussian [2], FSGS [3], CoR-GS [22], and Binocular3DGS [10]. Our two proposed strategies, dropout regularization and opacity noise injection, are applied to each method to evaluate their effectiveness. Ablation studies are primarily conducted on Binocular3DGS. We follow the official implementations for all baselines. On LLFF, our reproduced results for DNGaussian and CoR-GS slightly differ from the reported values. For Blender, DNGaussian uses different training settings for different scenes, while we adopt a unified setup for fair comparison, which leads to lower results. We also re-run Binocular3DGS with white backgrounds to align with other methods. Baseline results in Table 2 are reported as original / reproduced, showing the original paper values and our re-implementation under the unified setup.
Metrics. To evaluate the rendering quality under novel views, we compute standard image quality metrics including average PSNR (peak signal-to-noise ratio), SSIM [56], and LPIPS [57] on each dataset. To assess the impact of different strategies on co-adaptation, we additionally report Co-Adaptation Score (CA) (Section 3.2) computed on both training views and novel test views.
@@ 5 Conclusion
In this paper, we introduce the concept of co-adaptation in 3D Gaussian Splatting (3DGS) [1] and reveal its link to appearance artifacts under sparse-view settings. We propose the Co-Adaptation Score (CA) metric to quantify this effect and show that higher view density naturally reduces co-adaptation. Motivated by this, we propose two simple strategies—dropout regularization and opacity noise injection—that effectively mitigate co-adaptation and improve sparse-view novel view synthesis. We hope our findings inspire future work on mitigating co-adaptation in learned 3D representations.
@@ A.2 Extended Qualitative Results
We present additional qualitative results on the LLFF [47], DTU [51], and Blender [52] datasets to further illustrate the visual improvements brought by our co-adaptation suppression strategies.
LLFF. We first examine the LLFF dataset, where improvements are clearly observed both in rendered images and Gaussian point clouds. As shown in Figure 9, our methods yield point clouds that are cleaner and more geometrically coherent. The baseline 3D Gaussian Splatting (3DGS) [1] often produces scattered, colorful speckles, which are significantly suppressed after applying our co-adaptation regularization. These visual improvements in point clouds directly lead to better rendering in novel views. For instance, in the trex scene, fossil bones exhibit high-frequency flickering and speckling in the baseline, which are clearly alleviated with our methods. In the flower scene, petals and leaves display incorrect colorization under novel views, while the regularized versions better preserve the natural appearance. Similarly, in the fern scene, severe color inconsistencies appear on the pole, ground, and leaf textures under the baseline, but these are significantly reduced when co-adaptation is suppressed. Overall, the results suggest that reducing co-adaptation enhances both structural consistency and appearance fidelity in sparse-view 3DGS.
@@
Method Comparison. Figure 12 presents a visual comparison across multiple 3DGS methods, using Binocular3DGS [10] as the baseline. Binocular3DGS is a strong baseline with geometry-aware initialization, yet our co-adaptation suppression strategies (dropout and opacity noise) still lead to noticeable improvements. In the flower scene, existing sparse-view approaches exhibit occasional color speckles and geometry gaps; applying our methods further reduces these artifacts and yields more complete, coherent novel view reconstructions.
@@ A.3 Exploration of Noise Injection on Different Gaussian Attributes
where $s$ denotes the original scale and $\sigma$ is the noise strength. The results of this experiment are summarized in Table 6. Compared to the Binocular3DGS baseline, scale noise yields limited improvements in overall reconstruction metrics such as PSNR, SSIM [56], and LPIPS [57]. However, it reduces the co-adaptation strength (CA) on both training and novel views, indicating that scale perturbation helps decouple Gaussian interactions to some extent. This finding aligns with our visual observations in Figure 13, where scale noise improves scene-level geometric coherence in some cases, but introduces mild local blurring in object details. The observed trade-off can be explained by the functional role of scale in Gaussian rendering: perturbing scale slightly alters the spatial support of Gaussians, which may improve coverage and fusion in sparse-view settings, but risks oversmoothing fine structures. As a result, while scale noise contributes to better scene fusion in some examples, it reduces local detail sharpness, leading to modest overall gains.
@@ A.7 Comparison with Alternative Strategies for Reducing Co-Adaptation
  • Baseline w/ AbsGS [58]: We apply pixel-wise absolute gradient accumulation for Gaussian densification, replacing the standard gradient used in vanilla 3DGS. This method promotes the preservation of fine structures and reduces appearance artifacts, which may help alleviate view-specific co-adaptation.
  • Baseline w/o Opacity Decay [10]: We remove the default multiplicative decay applied to Gaussian opacity parameters ($\times 0.995$ per iteration). Without this decay, redundant Gaussian opacities tend to maintain, causing higher co-adaptation scores, reduced sparsity, and more pronounced floating artifacts.
@@
w/ AbsGS [58] 21.61 0.754 0.163 0.001446 0.001593
w/o opacity decay [10] 20.26 0.708 0.202 0.004413 0.004830